# Keyframer: Empowering Animation Design using Large Language Models

### v1

@@ 7. Discussion
Supporting design through iterative prompting. In our user study, the majority of participants took on a “decomposed” prompting strategy (explained in Section 6.1), with which they iteratively build up prompts, alternating between requests for animating new items and refinement of existing elements. This prompting strategy is a direct result of the + Add New Prompt feature in Keyframer as part of its affordances for supporting design iteration (explained in Section 4).
This iterative way of prompting LLMs can effectively support collaboration between animation creators and the LLM. As users may not initially have a holistic picture on what they want their animation to look like, this prompting strategy allows them to iteratively develop ideas in collaboration with the LLM, exploring what is possible through small, incremental steps. In addition, this prompting strategy aids the understandability and control of the LLM as incremental changes can help designers connect their prompts to changes in generated code. These findings echo recent studies on “prompt chaining” with LLMs (Wu et al., 2022), which refers to the mechanism by which the output of one step becomes the input for the next and can improve system transparency and controllability.
However, different from (Wu et al., 2022), we found that with Keyframer, users were adaptively and fluidly refining their goals through responding to the output from the model. This prompting strategy also contrasts with the one-shot prompting process commonly found in current LLM-based text-to-image systems, which encourage users to describe their goals in a detailed, single prompt and refine the same prompt to achieve a desired output (Kulkarni et al., 2023). We argue that the decomposed prompting mechanism supported by Keyframer is more human-centered, allowing users to engage in an iterative design process (Kelly, 2002) by gradually forming design goals as they converse with the LLM across multiple prompts. We hope that future LLM-powered design tools can consider this prompting strategy to increase user control and interpretability.
LLMs can empower a broad range of users to create animations in new ways. Several results positively support the use of LLMs to empower both beginners and professional animators. First, we found that the majority of prompts users created (84.4%) were semantic prompts descriptive of parameters like visual effects, timing, and coordination without using keywords commonly used in animation software. This result suggests that users do not need to have prior animation experience to effectively create animations using natural language.
Both “low code” (LC) and “high code” (HC) users engaged with editing LLM-generated code (80.0% and 71.4% respectively), indicating the value in enabling direct refinement of generated designs even for users without prior CSS experience. This is likely because many CSS properties can be interpretable to non-experts (such as duration and opacity). Showing a direct connection between rendered animations and the code behind it can support learning opportunities for creative computing. But even expert programmers appreciated the ability of the LLM to generate boilerplate code, saving them time to look up CSS syntax externally.
There were multiple occasions where the LLM generated designs that stimulated a user’s creativity, whether through unexpected design output or through users explicitly prompting for new ideas; these results indicate that LLMs can encourage creativity in the domain of animations.
Open UX questions for integrating design variants. While users who requested multiple designs often found this feature valuable for overcoming creative blocks or comparing alternatives, users did not necessarily choose to use this feature on their own. All but one of the 10 participants who generated multiple designs did so only in the second activity after the facilitator asked whether they intentionally did not use the feature in the first activity. While some participants forgot about this feature, potentially due to the lack of UI indicators for generating multiple designs, many users described how their typical workflows do not normally accommodate these design explorations, particularly when they have clear goals of what they want to create. While some systems like Midjourney automatically generate multiple designs, our participants expressed wanting to keep this feature opt-in, sharing how they would prefer as fast of a response as possible.
This time tradeoff, in which users have to actively decide whether asking for multiple designs is worth the longer response time, appears to impact use of this feature, along with parallel requests from participants for the LLM to produce richer, more unexpected design variants. We imagine multiple approaches to further investigate the use of variants, including fine tuning models to improve generated animation quality, considering alternative user interface for requesting variants (perhaps beyond or in conjunction with prompting), and providing feedback loops for users to understand what levers are available for better steering the LLM towards creating more distinct options.
Opportunities for improving interpretability. Participants were able to gain insight into the LLM’s interpretation of their prompts through the combination of inline animation previews, descriptive generated code (such as keyframe names like ‘flameFlicker,’ and domain-relevant CSS properties like ‘opacity’ and ‘fill’), and the generated explanations accompanying each design. Beyond these affordances, users offered several ideas for improving interpretability, many of which involve visual highlighting of relevant elements in both the SVG and generated animations. We heard requests for improving interpretability at three stages: during prompting, before design generation, and after the design has been generated. During prompting, participants suggested syntax highlighting when an identifier from the SVG was recognized in the prompt input field, helping them confirm the LLM’s interpretation of keywords in their prompt. Before generating a design, some participants wanted the LLM to proactively confirm its understanding by asking a series of questions such as if the animation should happen repeatedly or just once. Once a user request has been made, others saw the multiple variant feature as a means to confirm the range of ways an LLM might interpret their prompt.
@@ 8. Limitations and Future Work
A common user request shared by 7 of the 13 participants was to be able to edit the underlying illustration, which is not easily achieved with CSS alone. Not being able to edit the SVG itself prevents actions like adding new items to a scene, modifying the shape of objects, or re-grouping objects, which might limit a user’s creativity, as described by EP11: “There’s not a lot of space for brainstorming because a lot of decisions were made, like the shape of the rocket or the font.” However, while an earlier version of Keyframer enabled generating CSS and SVG edits, it led to much longer response times and likelihood of exceeding token limitations because of the amount of code in the SVG itself. So while it might be ideal for users to be able to leverage the LLM to make changes to the illustration metadata directly, doing so can impact the performance of the system.
We designed Keyframer to resemble the ChatGPT interface in which prompting happens sequentially and linearly. However, this linear representation was not necessarily representative of how some of our users approach animations. For example, EP12 stated that he would typically tune the movement of two items independently and simultaneously throughout the process of creating an animation. A non-linear representation may help users expecting to tweak and merge different animated elements together, as explored in related work (Angert et al., 2023; Dang et al., 2023).
Finally, many users wanted prompt-based interfaces alongside direct manipulation controls they are accustomed to in applications like After Effects. Users such as EP13 described how transform operations can be faster through direct manipulation (moving items to different places for individual keyframes), while more complex, non-translational animations might be easier through prompting. We imagine future AI-powered animation tools combining the affordances of direct manipulation with prompting for streamlining the definition and editing of animated properties.
@@ 9. Conclusion
Through Keyframer, we introduced how LLMs can shape future animation design tools by supporting iterative prototyping across exploration and refinement stages of the design process. Keyframer takes advantage of the code generation capabilities of LLMs and affordances for natural language input to empower users to generate animations using highly semantic prompts. Through our exploratory user study, we contribute a taxonomy of semantic prompting styles observed from users describing motion in natural language, ranging from those describing timing and coordination, along with more open-ended prompts soliciting ideas from the LLM (e.g., “make it look cool”). Further, we illustrated how we can enable animation creators to maintain creative control by providing pathways for iteration, with users alternating between prompting and editing generated animation code to refine their designs, and users building up designs through sequential prompting (what we describe as ‘decomposed’ prompting). Yet, we also found that Keyframer users found value in unexpected LLM output that helped spur their creativity. One way we observed this serendipity was with users testing out Keyframer’s feature to request design variants. Through this work, we hope to inspire future animation design tools that combine the powerful generative capabilities of LLMs to expedite design prototyping with dynamic editors that enable creators to maintain creative control in refining and iterating on their designs.
@@ A.1. Formative Interviews Participants
Our formative interview participants consisted of professionals who create animators as part of their work at a large technology company.
@@ A.2. Illustration Creation
A.2.1. SVG Pre-Processing
For our user study, we did not expect users to need to have any knowledge of SVG syntax to successfully complete our tasks — we wanted users to pay attention only to the identifiers for items in the scene (which could be used for prompting the LLM). All other SVG properties (like fills or path origins) could be ignored. Additionally, to reduce the LLM response time and likelihood of encountering token limitations (as the SVG itself is passed in each request to GPT-4), we took efforts to make the SVGs as small as possible by reducing the lines of code in the XML. To accomplish both of these goals, we pre-processed the SVGs through the following steps:
@@ A.2.1. SVG Pre-Processing
After pre-processing, the SVG for Activity 1 consisted of 42 lines of code, and the SVG for Activity 2 consisted of 40 lines of code, translating to 2,503 and 3,667 tokens respectively for GPT-4151515https://platform.openai.com/tokenizer.
@@ A.7. CSS Errors
Type Count
Example
Classname applied to keyframe 6 (40.0%)
.design-9 @keyframes planet-rotate {} instead of @keyframes planet-rotate {}
Class notation instead of id 3 (20.0%)
.design-0 .flame instead of .design-0 #flame
Incorrect class refactoring 2 (13.3%)
.design-0 #rocketship {} .design-3 #exhaust {} (both should use .design-3)
Typo 2 (13.3%)
¡stle¿ instead of ¡style¿
@@
Undefined variable 1 (8%)
calc(30deg * var(—random)) when random has not been defined

### v2

@@ 7. Discussion
We found that a key aspect of Keyframer’s design was the combined features of supporting semantic natural language prompts, along with direct editing of the rendered animations through code. Natural language input can support users in specifying their goals, and the majority of prompts users created (84.4%) were semantic prompts that describe parameters like visual effects, timing, and coordination without using keywords commonly used in animation software. Simultaneously, even “low code” users found that direct modification of the LLM-generated code and parameters was helpful for fine-tuning their designs. This means that in the context of using LLMs to generate code-based visual artifacts, tool developers should spend effort building direct manipulation interfaces and not simply fall back to a simpler chat-only design.
Our participants appreciated the ability to progressively construct their prompts, generating their animated creations over multiple iterations. This contrasts with the one-shot interfaces commonly found in text-to-image generators, where users must describe their goals in a detailed, single prompt (Kulkarni et al., 2023) and where image quality is often correlated with lengthy prompts (Xie et al., 2023). This was particularly enabled by the choice to save the generated artifacts at each prompt iteration, and to allow users to go back and create a new prompt starting with any previous iteration. Not only did this implementation choice support progressive iteration, but it also enabled the “decomposed” prompting strategy that was used by the majority of our participants. This ability to construct from scratch and adapt design goals over time is in contrast to related animation work that centers on automated approaches to generating animations (Liu et al., 2024; Ma and Agrawala, 2025), which may run the risk of design fixation in response to seemingly-fully-realized designs (Anderson et al., 2024; Suh et al., 2024).
Explicit support for multiple variations also assists in the iterative process and is another feature of Keyframer that should be included where possible in future animation tools. We found this feature was most useful in the exploratory design phase, but could be a hindrance during fine-tuning later in the design process. Much work remains to build even better tools for generating and displaying multiple variations. A particularly fruitful area for future work might be to explore methods for determining the best number of variations to generate based on where the user is in the design process, ensuring that those variations are clearly distinct, and finding better ways to visualize differences between variations when they are subtle.
There remain several open questions about how LLMs might be effectively integrated into animation design processes.
Enabling user feedback on model interpretability In the context of applying our categorization of semantic prompts, we imagine opportunities for systems to provide enhanced user feedback on model interpretability of users prompts. For example, during prompting, a system might proactively display its understanding by both highlighting keywords in the user prompt along with elements in the illustration, which may help users confirm how LLMs are matching their prompt keywords with individual components within a graphic. Before generating a design, some users in our study suggested having the LLM ask questions to confirm its understanding, which may assist users with clarifying their intention through revising their wording. After generation, features like Keyframer’s ability to display variants could be used to expose the range of ways an LLM might interpret their prompt, enabling users to choose from several options the one that best fits their vision.
Supporting task decomposition for non-linear design processes While Keyframer was designed to resemble the ChatGPT interface in which prompting happens sequentially and linearly, this linear representation was not necessarily representative of how some of our users approach animations. For example, EP12 stated that he would typically tune the movement of two items independently and simultaneously throughout the process of creating an animation. Future work integrating non-linear representation may help users expecting to tweak and merge different animated elements together, as explored in related work (Angert et al., 2023; Dang et al., 2023). Interfaces that visually represent task decomposition on individual elements within a scene may enable more granular and flexible refinement.
Enabling edits on underlying graphics A common user request was to edit the underlying illustration using the LLM, which is not easily achieved with CSS alone. Not being able to edit the SVG itself prevents actions like adding new items to a scene, modifying the shape of objects, or re-grouping objects. While an earlier version of Keyframer enabled generating CSS and SVG edits, it led to much longer response times and likelihood of exceeding token limitations because of the amount of code in the SVG itself. Future work can consider how LLMs can be used to edit both animation properties and the illustration itself to offer even more creative opportunities.
Mixing prompting and direct manipulation Many animators are used to direct manipulation via timeline-based interfaces such as those available in After Effects. For example, EP13 described how transform operations can be faster through direct manipulation, while more complex, non-translational animations might be easier through prompting. An open question is how to design future AI-powered animation tools combining the affordances of direct manipulation (such as that offered with (Masson et al., 2024)) with prompting for streamlining the definition and editing of animated properties.
Designing animations with interactivity Finally, in Keyframer we explored the creation of animations for non-interactive contexts where the animation always plays exactly the same way every time. Other animation tasks, such as designing how a button appears when pressed or how a user interface responds to a gesture, need to take the user’s interaction into account during playback. This type of task is more complicated, as it requires anticipating possible interactions and viewing previews of diverse interactions that affect the animation design in different ways. LLMs might serve an additional purpose in such a tool by helping both with design as in Keyframer, but also in helping the designer understand possible interactions and identify edge cases in their designs.
@@ 8. Limitations and Future Work
There are several limitations in what Keyframer supports that could be expanded in future work. First, a common user request was to be able to edit the underlying illustration, which is not easily achieved with CSS alone. Not being able to edit the SVG prevents actions like adding new items to a scene or modifying the shape of objects, which may limit a user’s creativity. However, while an earlier version of Keyframer enabled generating CSS and SVG edits, it led to much longer response times and likelihood of exceeding token limitations because of the amount of code in the SVG itself. So while it might be ideal for users to leverage the LLM to make changes to illustration metadata directly, doing so can impact system performance.
Combining direct manipulation with prompting, such as described in (Masson et al., 2024) and Huang et al. (2024), may streamline the definition and editing of animated properties. For example, EP13 described how transforms may be faster through direct manipulation, while more complex animations might be easier through prompting. Further, future work could translate text-based CSS animation code to a visual timelines that may better coordination of timing between elements, such as that used in Chrome’s Animations panel UI. [^18] https://developer.chrome.com/docs/devtools/css/animations/#animations-ui
We utilized GPT-4 as our underlying LLM because of its availability and performance at the time of our experiments, but new models may yield different prompting approaches. In particular, multimodal models are now generally available, but were not at the time of this work. We used SVG as our image format because of its text-driven nature and its feature of labeling layers with natural language strings, which worked well with the text-only LLM that we used in Keyframer. Newer systems built on multimodal models could have far greater flexibility.
Finally, further work is needed to understand prompting practices for more complex illustrations, such as those used in UI animations and transitions. Yet, we believe that animated 2D illustrations such as those used in our user studies are representative of many potential applications like animated album art, emojis, or user avatars.
@@ 9. Conclusion
Through Keyframer, we envisioned how future animation design tools could be designed to support highly semantic natural language prompting, leverage and preview the outcomes of code generation, and enable design iteration through direct editing.
Through our user study, we contribute a categorization of semantic prompting styles observed from users describing motion in natural language. We revealed ways to provide pathways for iteration, where users can alternate between prompting and editing generated animation code to refine their designs. We also unpacked how animation creators can gradually build up and explore ideas through sequential ‘decomposed’ prompting. Through this work, we hope to inspire future animation tools that leverage the powerful generative capabilities of LLMs to expedite design prototyping, with dynamic editors that enable creators to maintain creative control throughout their design process.
@@ A.1. Formative Interviews Participants
Our formative interview participants consisted of professionals who create animators as part of their work at a large technology company, as displayed in Table 3.
@@ A.2. Illustration Creation
SVG Pre-Processing. For our user study, we did not expect users to need to have any knowledge of SVG syntax to successfully complete our tasks — we wanted users to pay attention only to the identifiers for items in the scene (which could be used for prompting the LLM). All other SVG properties (like fills or path origins) could be ignored. Additionally, to reduce the LLM response time and likelihood of encountering token limitations (as the SVG itself is passed in each request to GPT-4), we took efforts to make the SVGs as small as possible by reducing the lines of code in the XML. To accomplish both of these goals, we pre-processed the SVGs through the following steps:
@@
After pre-processing, the SVG for Activity 1 consisted of 42 lines of code, and the SVG for Activity 2 consisted of 40 lines of code, translating to 2,503 and 3,667 tokens respectively for GPT-4191919https://platform.openai.com/tokenizer.
@@ A.5. Full Prompt for Generating CSS Animation Code with an Input SVG
Note that while the prompt asks the LLM to retain existing CSS and only append new CSS below, in practice we found that the LLM often modifies existing code when it is more efficient. For example, if a user prompts the system to increase the speed for a particular animated feature, the LLM updates the ‘animation-duration’ property of the existing CSS code, rather than adding an override for the timing below it.
@@ A.7. CSS Errors
Type Count Example
Classname applied to keyframe 6 (40.0%) .design-9 @keyframes planet-rotate {} instead of @keyframes planet-rotate {}
Class notation instead of id 3 (20.0%) .design-0 .flame instead of .design-0 #flame
Incorrect class refactoring 2 (13.3%) .design-0 #rocketship {} .design-3 #exhaust {} (both should use .design-3)
Typo 2 (13.3%) ¡stle¿ instead of ¡style¿
@@
Undefined variable 1 (8%) calc(30deg * var(—random)) when random has not been defined